University of Ottawa NMR Facility Web Site

Please feel free to make suggestions for future posts by emailing Glenn Facey.



Sunday, September 9, 2007

All Spectrometers Disabled After Unexpected Power Fluctuation

On Friday evening September 7th, there was a power fluctuation at the university which disabled all 6 of our NMR spectrometers. The instruments were all restarted on Saturday between 11:30 am and 1:30 pm. The NMR experiments on the AVANCE 400 which had errors were resubmitted to the queue. I restarted the experiments that were running on the ASX 200 and AVANCE 500. The student who was running the INOVA 500 should contact Cheryl. There did not seem to be any user of the AVANCE 300.

1 comment:

Anonymous said...

Hey Glenn - thanks for restarting the Hg-199 experiment. I suspect I might have to re-tweak the match but we'll see...

D.